# Not All Language Model Features Are Linear

### v1

@@ Appendix D Toy Case of Training SAEs on Circles
To explore how SAEs behave when reconstructing irreducible features of dimension $d_{f}>1$, we perform experiments with the following toy setup. Inspired by the circular representations of integers that networks learn when trained on modular addition [35, 27], we create synthetic datasets of activations containing multiple features which are each 2d irreducible circles.
@@ Appendix E Training Mistral SAEs
Our Mistral 7B [21] sparse autoencoders (SAEs) are trained on over one billion tokens from a subset of the Pile [12] and Alpaca [39] datasets. We train our SAEs on layers 8, 16, and 24 out of 32 total layers to maximize coverage of the model’s representations. We use a 16×16$\times$16 × expansion factor, yielding a total of 65536 dictionary elements for each SAE.
To train our SAEs, we use an $L_{p}$ sparsity penalty for $p=1/2$ with sparsity coefficient $\lambda=0.012$. Before an SAE forward pass, we normalize our activation vectors to have norm $\sqrt{d_{model}}=64$ in the case of Mistral. We do not apply a pre-encoder bias. We use an AdamW optimizer with weight decay $10^{-3}$ and learning rate 0.0002 with a linear warm up. We apply dead feature resampling [6] five times over the course of training to converge on SAEs with around 1000 dead features.
@@ F.2 Mistral 7B methods and results
As future work, we think it would be exciting to develop better clustering techniques for SAE features. Our graph based clustering technique could likely be improved by more recent efficient and high-quality graph based clustering techniques, e.g. hierarchical agglomerate clustering with single-linkage [25]. Additionally, we believe we would see a large improvement by setting edge weights to be a combination of both the cosine and Jaccard similarity of the dictionary elements, e.g. max(cosine, Jaccard).
@@ G.1 Assets Information
We use the following open source models for our experiments: Llama 3 8B [2] (custom Llama 3 license https://llama.meta.com/llama3/license/), Mistral 7B [21] (released under the Apache 2 License), and GPT-2 [41] (modified MIT license, see https://github.com/openai/gpt-2/blob/master/LICENSE).
@@ G.2 Machine Information
Intervention experiments were run on two V100 GPUs using less than $64$ GB of CPU RAM; all experiments can be reproduced from our open source repository in less than a day with this configuration. We use the TransformerLens library [34] for intervention experiments. $\epsilon$-mixture index measurements on toy datasets took about one minute each, on 8GB of CPU RAM. EVR experiments take seconds on 8GB of CPU RAM and are dominated by time taken to human-interpret the RGB plots.
@@ Appendix H More Weekdays and Months Plots and Details
For simplicity of presentation, we clip all of the (few) negative intervention averages to 00 (prior work [48] has also found negative-effect attention heads during patching experiments).

### v2

@@ Appendix D Toy Case of Training SAEs on Circles
To explore how SAEs behave when reconstructing irreducible features of dimension $d_{f}>1$, we perform experiments with the following toy setup. Inspired by the circular representations of integers that networks learn when trained on modular addition (Nanda et al., 2023a; Liu et al., 2022), we create synthetic datasets of activations containing multiple features which are each 2d irreducible circles.
@@ Appendix E Training Mistral SAEs
Our Mistral 7B (Jiang et al., 2023) sparse autoencoders (SAEs) are trained on over one billion tokens from a subset of the Pile (Gao et al., 2020) and Alpaca (Peng et al., 2023) datasets. We train our SAEs on layers 8, 16, and 24 out of 32 total layers to maximize coverage of the model’s representations. We use a 16×16$\times$16 × expansion factor, yielding a total of 65536 dictionary elements for each SAE.
To train our SAEs, we use an $L_{p}$ sparsity penalty for $p=1/2$ with sparsity coefficient $\lambda=0.012$. Before an SAE forward pass, we normalize our activation vectors to have norm $\sqrt{d_{model}}=64$ in the case of Mistral. We do not apply a pre-encoder bias. We use an AdamW optimizer with weight decay $10^{-3}$ and learning rate 0.0002 with a linear warm up. We apply dead feature resampling (Bricken et al., 2023) five times over the course of training to converge on SAEs with around 1000 dead features.
@@ F.2 Mistral 7B methods and results
As future work, we think it would be exciting to develop better clustering techniques for SAE features. Our graph based clustering technique could likely be improved by more recent efficient and high-quality graph based clustering techniques, e.g. hierarchical agglomerate clustering with single-linkage (Lattanzi et al., 2020). Additionally, we believe we would see a large improvement by setting edge weights to be a combination of both the cosine and Jaccard similarity of the dictionary elements, e.g. max(cosine, Jaccard).
@@ G.1 Assets Information
We use the following open source models for our experiments: Llama 3 8B (AI@Meta, 2024) (custom Llama 3 license https://llama.meta.com/llama3/license/), Mistral 7B (Jiang et al., 2023) (released under the Apache 2 License), and GPT-2 (Radford et al., 2019) (modified MIT license, see https://github.com/openai/gpt-2/blob/master/LICENSE).
@@ G.2 Machine Information
Intervention experiments were run on two V100 GPUs using less than $64$ GB of CPU RAM; all experiments can be reproduced from our open source repository in less than a day with this configuration. We use the TransformerLens library (Nanda & Bloom, 2022) for intervention experiments. $\epsilon$-mixture index measurements on toy datasets took about one minute each, on 8GB of CPU RAM. EVR experiments take seconds on 8GB of CPU RAM and are dominated by time taken to human-interpret the RGB plots.
@@ Appendix H More Weekdays and Months Plots and Details
Figure 19: Layer $30$ Mistral 7B activations for [morning/evening] on [Monday/Tuesday/…/Sunday], plotted projected into the PCA plane for [Monday/Tuesday/…/Sunday].
@@
For simplicity of presentation, we clip all of the (few) negative intervention averages to 00 (prior work (Zhang & Nanda, 2023) has also found negative-effect attention heads during patching experiments).
Finally, in Fig. 19, we show another example of the continuity of the circular days of the week representation in Mistral 7B.
@@ Appendix I Patching
To isolate the rough circuit for Weekdays and Months, we perform layer-wise activation patching on $40$ random pairs of prompts. The results, displayed in Fig. 18 show that the circuit to compute $\gamma$ consists of MLPs on top of the $\alpha$ and $\beta$ tokens, a copy to the token before $\gamma$, and further MLPs there (roughly similar to prior work studying arithmetic circuits (Stolfo et al., 2023)). Moreover, fine-grained patching in Appendix J shows that there are just a few responsible attention heads for the writes to the token before $\gamma$. However, patching alone cannot tell use how or where $\gamma$ is represented. For that, we need a new technique, which we expand on in the next section.
Appendix J Explanation via Regression (EVR)
In this section, we introduce a new technique for empirically explaining hidden representations in algorithmic problems: explanation via regression (EVR). Given a set of tokens $T$ with a corresponding set of hidden states $X_{i,l}$, we explain the variance in $X_{i,l}$ by adding together hand-chosen functions of $t$. This gives us an explanation of what the transformation $T\rightarrow X_{i,l}$ computes. For a given choice of explanation functions $\{{\mathbf{g}}_{i}(t)\}$, the $r^{2}$ value of a linear regression from $\{{\mathbf{g}}_{i}(t)\}$ to $X_{i,l}$ gives a measure of the explained variance in $X_{i,l}$. But what functions ${\mathbf{g}}_{i}$ should we choose?
In  Fig. 20, we perform EVR on the layer 17-29 hidden states of Mistral 7B on the Weekdays task; additional deconstructions are in Appendix J. We find that a circle in $\gamma$ develops and grows in explanatory power; we plot the layer $25$ residuals after explaining with one hot functions in $\alpha$ and $\beta$ (i.e. ${\mathbf{g}}_{1}=[\alpha=0],{\mathbf{g}}_{2}=[\beta=1],{\mathbf{g}}_{3}=[\alpha=1],\ldots$) in Fig. 8 to show this incredibly clear circle in $\gamma$. This suggests that the models may be generating $\gamma$ by using a trigonometry based algorithm like the “clock” (Nanda et al., 2023a) or “pizza” (Zhong et al., 2024) algorithm in late MLP layers.